I went to get a sandwich around 10:20 AM and when I came back BS was moving like crazy. We shutdown the watchdog.
We look at the spectra of the OSEMs (attachment 1). Clearly, the UR sensing is bad.
We took the BS sattelite box out. Anchal opened the box and nothing seemed wrong visually. We returned the box and connected it to the fake OSEM box. The sensor spectra seemed normal.
We connected the box to the vacuum chamber and the spectra is still normal (attachment 2).
We turn on the coils and the motion got damped very quickly (RMS <0.5mV).
Either the problem was solved by disconnecting and connecting the cables or it will come back to haunt us.